# Universal Frequency Correlations and Recurrence Statistics of Complex Impedance Matrices

### v1

@@ 4 Discussion
While we do indeed see a $\frac{1}{\epsilon}$ dependence for all the curves in each panel, there is some vertical offset between the curves from different ensembles, illustrated by the color scale corresponding to the ensemble $\alpha$ value. We have determined two possible reasons for the spread out nature of the curves. When looking at the experimental data in panels (a) and (b), it can be seen that the red curves corresponding to ensembles with $\alpha\geq 5$ are degenerate, and only the blue, green and yellow curves fail to align with the scaling curve. This suggests that for small absorption there is some leftover $\alpha$ dependence to $\Lambda(Z^{\text{peak}})$ that is not properly captured by Eq. 6, which assumes high loss ($\alpha\gg 1$). However, the RMT numerical simulations in panels (c) and (d) result in $\Lambda(Z^{\text{peak}})/(\alpha^{1/2}\Delta)$ curves that seem degenerate for all $\alpha$ values. This implies the issue may not be in the theory but actually small inaccuracies in the $\Delta$ value used. Mode spacing in the RMT simulation is very well described by the prescribed $\Delta$, while for the experimental ensembles the value of $\Delta$ is just an approximation calculated using Weyl’s formula.
@@ 5 Conclusions
In this paper, we have presented an experimental study of the statistics of the RCM normalized impedance of microwave cavities. We have examined the two point frequency correlation functions, which agree nicely with the theoretical predictions. We also considered the question of how far on average one must look in frequency before seeing a repeated value of impedance. We demonstrated that this frequency interval $\Lambda$ depends on four parameters: the mean mode spacing $\Delta$, the absorption $\alpha$, the value of target impedance $Z^{\text{t}}$, and the tolerance $\epsilon$ on the desired impedance. Experimentally, $\Delta$ and $\alpha$ are inversely correlated and it can be difficult to tune one without affecting the other, which means it cannot be predicted whether $\Lambda$ will shrink or grow based on changes in just one of these parameters. Further, the PDF of $Z$ depends on $\alpha$, which introduces another, more complicated, dependence on loss to $\Lambda$. Using Random Matrix Theory simulations, we see statistical results that agree well with our data. The agreement between the experiment and RMT strongly imply our results are applicable to any sufficiently complicated wave scattering system, including those seen in optics or acoustics. An extension of this work is to consider not just special values of individual impedance elements, but special conditions on the entire $Z$ matrix, such as impedance matrix Exceptional Points, which are equivalent to scattering matrix Exceptional Points [39].
6 Appendix: Random Coupling Model Loss Parameter Determination
@@
Because this process uses an ensemble that involves both cavity manipulation and a range of frequencies, we are in a sense averaging the contributions of multiple $\alpha$ values. This is reasonable when $\alpha$ is large and the contributions aren’t too different from each other [72]. However, since the PDFs of $Z$ have sensitivity of shape and variance to $\alpha$ that scale roughly with the inverse of $\alpha$, it can be difficult to do the same when $\alpha<1$.
Our experimental systems with their in situ tunable perturbers allow us to make very high quality ensembles for which we can apply this process even in cases of very low loss. See the Supplemental Materials of Ref. [37] for an example where we simultaneously match all $8$ impedance PDFs of an $M=2$ channel graph to the RCM predicted PDFs with $\alpha=0.18$. This method is even applicable in the case of ultra low loss enabled through cryogenic environments, as shown in Ref. [73].

### v2

@@ 4 Discussion
While we do indeed see a $\frac{1}{\epsilon}$ dependence for all the curves in each panel, there is some vertical offset between the curves from different ensembles, illustrated by the color scale corresponding to the ensemble $\alpha$ value. We have determined two possible reasons for the spread out nature of the curves. When looking at the experimental data in panels (a) and (b), it can be seen that the red curves corresponding to ensembles with $\alpha\geq 5$ are degenerate, and only the blue, green and yellow curves fail to align with the scaling curve. This suggests that for small absorption there is some leftover $\alpha$ dependence to $\Lambda(Z^{\text{peak}})$ that is not properly captured by Eq. 6, which assumes high loss ($\alpha\gg 1$). However, the RMT numerical simulations in panels (c) and (d) result in $\Lambda(Z^{\text{peak}})/(\alpha^{1/2}\Delta)$ curves that seem degenerate for all $\alpha$ values. This implies the issue may not be in the theory but actually small inaccuracies in the $\Delta$ value used. Mode spacing in the RMT simulation is very well described by the prescribed $\Delta$, while for the experimental ensembles the value of $\Delta$ is just an approximation calculated using Weyl’s formula. As for why the theory describes the statistical results of even low loss ensembles, both experimental and from RMT simulation, it is likely because the correlation functions remain qualitatively similar across all $\alpha$ values. Hence the simple geometrical argument for $\Lambda(Z^{\text{t}}=Z^{\text{peak}})=\frac{\pi\gamma R^{\text{RMS}}}{\epsilon}$ can be expected to hold.
@@ 5 Conclusions
In this paper, we have presented an experimental study of the statistics of the RCM normalized impedance of microwave cavities. We have examined the two point frequency correlation functions, which agree nicely with the theoretical predictions. We also considered the question of how far on average one must look in frequency before seeing a repeated value of impedance. We demonstrated that this frequency interval $\Lambda$ depends on four parameters: the mean mode spacing $\Delta$, the absorption $\alpha$, the value of target impedance $Z^{\text{t}}$, and the tolerance $\epsilon$ on the desired impedance. Experimentally, the mode spacing $\Delta$ and the mode overlap $\alpha$ are inversely correlated and it can be difficult to tune one without affecting the other, which means it cannot be predicted whether $\Lambda$ will shrink or grow based on changes in just one of these parameters. Further, the PDF of $Z$ depends on $\alpha$, which introduces another, more complicated, dependence on loss to $\Lambda$. Using Random Matrix Theory simulations, we see statistical results that agree well with our data. The agreement between the experiment and RMT strongly imply our results are applicable to any sufficiently complicated wave scattering system, including those seen in optics or acoustics. An extension of this work is to consider not just special values of individual impedance elements, but special conditions on the entire $Z$ matrix, such as impedance matrix Exceptional Points, which are equivalent to scattering matrix Exceptional Points.
6 Appendix A: Table of Parameters
Parameter Symbol Definition Relation to other Parameters
$Z^{\text{raw}}$ Raw impedance matrix Eq. 8
$\langle Z^{\text{raw}}\rangle$ Ensemble average of raw impedance. For a high quality ensemble with many realizations, becomes nearly identical to radiation impedance with small differences caused by short orbits. N/A
$Z$ Universal fluctuating impedance matrix, normalized to not contain system specific features Eq. 9
$Z^{\text{RCM}}$ Universal fluctuating impedance matrix as predicted by the Random Coupling Model and RMT Eq. 10
$Z^{\text{peak}}$ Impedance value that corresponds to the largest value of $\mathcal{P}(Z)$ $\mathcal{P}(Z^{\text{peak}})=\text{max}[\mathcal{P}(Z)]$
$Z^{\text{t}}$ Target impedance value of interest N/A
$\epsilon$ Tolerance on target impedance $Z^{\text{t}}$ N/A
$R^{\text{RMS}}$ Root mean square fluctuation of the normalized impedance $R^{\text{RMS}}=\sqrt{\langle(R_{pp^{\prime}})^{2}\rangle}$ and at large loss $R^{\text{RMS}}=\sqrt{\frac{1}{\pi\alpha}}$
$\alpha$ RCM loss parameter, quantifies degree of mode overlap N/A
$\Delta$ Mean frequency spacing of the modes N/A
$\gamma$ Mean spectral $Q$-width of the modes $\gamma=2\alpha\Delta$
$\delta$ Difference between two frequency values $f_{a}$ and $f_{b}$ $\delta=|f_{a}-f_{b}|$
$\Lambda$ Average frequency interval between repeated occurances of target impedance value $Z^{\text{t}}$ Eqs. 5-6
Table 1: Table of important parameters used in this work. The symbol used is given in the first column and the definition in the second column. How that parameter is related to other parameters is given in the last column.
7 Appendix B: Random Coupling Model Loss Parameter Determination
@@
Because this process uses an ensemble that involves both cavity manipulation and a range of frequencies, we are in a sense averaging the contributions of multiple $\alpha$ values. This is reasonable when $\alpha$ is large and the contributions aren’t too different from each other [Fu2017]. However, since the PDFs of $Z$ have sensitivity of shape and variance to $\alpha$ that scale roughly with the inverse of $\alpha$, it can be difficult to do the same when $\alpha<1$.
Our experimental systems with their in situ tunable perturbers allow us to make very high quality ensembles for which we can apply this process even in cases of very low loss. See the Supplemental Materials of Ref. [Shaibe2025_CTD] for an example where we simultaneously match all $8$ impedance PDFs of an $M=2$ channel graph to the RCM predicted PDFs with $\alpha=0.18$. This method is even applicable in the case of ultra low loss enabled through cryogenic environments, as shown in Ref. [Yeh2013Cyro].